OpenStudy (anonymous):

what is the derivative of f(x)= 3cos^2 (x)

OpenStudy (raden):

chain rule is works f(x) = 3cos^2 (x) f ' (x) = 3 * 2 cos(x) * (-sin(x) f ' (x) = -3 * 2 cos(x) * sin(x) if you want simplify again, use the identity 2 cos(x) * sin(x) = sin(2x) so, f ' (x) = -3sin(2x)

OpenStudy (anonymous):

how did you get 2 cos(x) * (-sin(x)???

OpenStudy (raden):

derivative of (..)^2 = 2(...) derivative of cos(x) is -sin(x) that's called the chain rule

OpenStudy (anonymous):

is there a rule for that?

OpenStudy (raden):

yes, of course but you can use other methode if you want

OpenStudy (anonymous):

could you please show me?

OpenStudy (raden):

just use the identity : cos^2 x = (1+cos2x)/2

OpenStudy (raden):

f(x)= 3cos^2 (x) f(x)= 3(1+cos2x)/2 f(x) = 3/2(1+cos2x) f(x) = 3/2 + 3/2 cos2x now derive above

OpenStudy (raden):

derivative of 3/2 is 0 derivative of 3/2 cos(2x) is 3/2 * -2sin(2x) = -3sin(2x) so, f ' = 0 + -3sin(2x) = -3sin(2x)

OpenStudy (raden):

got it ?

OpenStudy (anonymous):

ill better have to write these down so that i can get a better picture

OpenStudy (anonymous):

thanks a lot. i got it now

OpenStudy (raden):

welcome